​Kaimalino Properties is evaluating a real estate investment of Mountain Ridge Estates. Management plans to buy the property today and sell it 10 years from today. The initial cost of the property is $18 million and the expected sale price is $56 million. What is the IRR of the investment? Enter your answer as a percentage and rounded to 2 DECIMAL PLACES. Do not include the percentage sign in your answer. ​

Enter your response below
Answer #1

Hi

We will find the IRR based on compounding interest formula

Here current price P=$18 million

Selling Price F=$56 million

Time t= 10 years

As per compounding interest formula

F= P*(1+IRR)^ t

56 = 18*(1+IRR)^10

(1+IRR)^10 = 56/18

1+IRR =(56/18)^(1/10)

1+IRR = 1.12019

IRR= 12.02%

Thanks